OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. G(4, 6), H(1, 5) Help please! ^_^

OpenStudy (anonymous):

*Standrard form* Ax+By=C where A,B,C are all integers and no fractions and A is not negative 1. Find the slope which is y2-y1/x2-x1 (4,6)=x1,y1 and (1,5)=x2,y2 2. 5-6/1-4= -1/-3= (1/3) the slope 3.Now plug this, along with either of the points into slope intercept form which is y=mx+b where m is the slope(1/3) and b is the y int, the last part we need for the standard form 4. We'll just use (1,5) 5=1/3(1)+b 5=1/3+b 4+(2/3)=b so the slope intercept form of the equation is y=(1/3)x+4+(2/3) (1/3)x+y=4+(2/3) multiply this all by 3 to get rid of all the fractions to get x+3y=14

OpenStudy (anonymous):

Equation of a line passing through two points (x1,y1) and (x2,y2) is y−y1=[(y2−y1)/(x2−x1)](x−x1) y-6 = [(5-6)/(1-4)](x-4) =1/3(x-4) =1/3x-4/3 =(1/3)x+(14/3)(ans) when i was in 5 th grade, i had face this co ordinate geometry : http://www.mathcaptain.com/geometry/coordinate-geometry.html problem but thanks for math teacher for helping me out.